13, what is the o nacasso
Answer #1

Given:

Ka = 1.7*10^-5

use:

pKa = -log Ka

= -log (1.7*10^-5)

= 4.7696

use:

pH = pKa + log {[conjugate base]/[acid]}

4.44 = 4.77+log {[CH3CO2-]/[CH3CO2H]}

log {[CH3CO2-]/[CH3CO2H]} = -0.3296

[CH3CO2-]/[CH3CO2H] = 0.4682

Answer: 0.468